### v1

@@ 7 Conclusion and future outlook
In this section, we consider further details of the conservative, FV semi-Lagrangian method outlined in Section 5.3. Section SM3.1 presents relevant details for how the numerical flux is implemented, Section SM3.2 develops a truncation error estimate, and then Section SM3.3 describes how this estimate is used to develop the expression for the associated truncation error operator ${\cal T}_{\rm direct}^{n}$ used in the coarse-grid operator (58).
SM3.1 Further details on numerical flux
@@
$\widetilde{f}_{i+1/2}=\int_{\widetilde{\xi}_{i+1/2}}^{\widetilde{x}_{i+1/2}}R_{i+1/2}[\bar{\bm{e}}^{n}](x)\operatorname{d}\!x+\int_{\widetilde{x}_{i+1/2}}^{x_{i+1/2}}R_{i+1/2}[\bar{\bm{e}}^{n}](x)\operatorname{d}\!x,$ (SM10)
where, recall, $\widetilde{x}_{i+1/2}$ is the FV cell interface immediately to the east of the departure point $\widetilde{\xi}_{i+1/2}$; see Fig. 1. The second integral in (SM10) is evaluated exactly, noting that it is simply a sum of cell averages across the cells in the interval $x\in[\widetilde{x}_{i+1/2},x_{i+1/2}]$. The first integral in (SM10) is more complicated. Recall that $R_{i+1/2}[\bar{\bm{e}}^{n}](x)$ is a polynomial reconstruction of $e^{n}(x)$ and it is constructed by enforcing that it preserves cell averages of $e^{n}(x)$ across the FV cells used in its stencil. For the remainder of this discussion, let us drop temporal subscripts since they are not important.
@@
Recall from Section 5.3 that we consider semi-Lagrangian methods of order $p=2k-1$ in space, with $k=1,2,\ldots$. To this end, let $R_{i+1/2}[\bar{\bm{e}}^{n}](x)$ be a reconstruction polynomial using a stencil consisting of $p=2k-1$ cells. Further, suppose the stencil is centered about the cell containing the departure point. For simplicity of notation, suppose that the cell containing the departure point, i.e., $x\in[\widetilde{x}_{i+1/2}-h,\widetilde{x}_{i+1/2}]$, is cell ${\cal I}_{0}$. Then, the cells in the stencil are $\{{\cal I}_{-(k-1)},\ldots,{\cal I}_{k-1}\}$. [^7] In the notation of Section SM1, the stencil has a total number of cells $k\mapsto 2k-1=p$, and a left shift of $\ell=k-1$. Define $E(x)$ as the primitive of $e(x)$, i.e., $E(x)=\int_{-\infty}^{x}e(\zeta)\operatorname{d}\!\zeta$ and $E^{\prime}(x)=e(x)$. Then, based on Section SM1, we know that $R_{i+1/2}[\bar{\bm{e}}](x)$ is a polynomial of degree at most $p-1$, and that it is the derivative of an interpolating polynomial $Q_{p}(x)$ of degree at most $p$,
$R_{i+1/2}[\bar{\bm{e}}](x)\equiv Q_{p}^{\prime}(x),$ (SM11)
@@
$Q_{p}(x_{j\pm 1/2})=E(x_{j\pm 1/2}),\quad j=-(k-1),\ldots,k-1.$ (SM12)
SM3.2 Error estimate
The estimate we develop herein and its proof share close parallels with the error estimate we developed previously in [14, Lemma 3.1] for a non-conservative, FD semi-Lagrangian method; however, the estimate itself and its proof are fundamentally different due to the fact that we are now considering a semi-Lagrangian method that is both conservative and based on finite volumes. The full estimate is split across Lemmas SM3.1 and SM3.3. Note that the proof of Lemma SM3.3 in particular requires to have read Section SM3.1.
Lemma SM3.1 (Error estimate for semi-Lagrangian method).
@@
$\tau^{n}_{i}$ $:=\Big{(}\bar{\bm{e}}(t_{n+1})-{\cal S}^{t_{n},\delta t}_{p,\infty}\bar{\bm{e}}(t_{n})\Big{)}_{i}=\frac{\widetilde{E}_{i+1/2}^{n}-\widetilde{E}_{i-1/2}^{n}}{h},$ (SM13)
@@
$\widetilde{E}_{i+1/2}^{n}$ $:=\int_{\widetilde{\xi}_{i+1/2}}^{\widetilde{x}_{i+1/2}}R_{i+1/2}[\bar{\bm{e}}(t_{n})](x)-e(x,t_{n})\operatorname{d}\!x.$ (SM14)
Proof SM3.2.
@@
$\tau^{n}_{i}$ $=\bar{e}_{i}(t_{n+1})-\bar{e}_{i}(t_{n})+\frac{\int_{\widetilde{\xi}_{i+1/2}}^{x_{i+1/2}}R_{i+1/2}[\bar{\bm{e}}(t_{n})](x)\operatorname{d}\!x-\int_{\widetilde{\xi}_{i-1/2}}^{x_{i-1/2}}R_{i-1/2}[\bar{\bm{e}}(t_{n})](x)\operatorname{d}\!x}{h}.$ (SM15)
@@
$\begin{split}&\tau^{n}_{i}=\left[\bar{e}_{i}(t_{n+1})-\frac{\int_{\widetilde{\xi}_{i-1/2}}^{x_{i-1/2}}e(x,t_{n})\operatorname{d}\!x+\int_{x_{i-1/2}}^{x_{i+1/2}}e(x,t_{n})\operatorname{d}\!x+\int_{x_{i+1/2}}^{\widetilde{\xi}_{i+1/2}}e(x,t_{n})\operatorname{d}\!x}{h}\right]+\\ &\frac{\Big{(}\int_{\widetilde{\xi}_{i+1/2}}^{x_{i+1/2}}R_{i+1/2}[\bar{\bm{e}}(t_{n})](x)-e(x,t_{n})\operatorname{d}\!x\Big{)}-\Big{(}\int_{\widetilde{\xi}_{i-1/2}}^{x_{i-1/2}}R_{i-1/2}[\bar{\bm{e}}(t_{n})](x)-e(x,t_{n})\operatorname{d}\!x\Big{)}}{h}.\end{split}$ (SM16)
The term inside the closed parentheses vanishes due to the local conservation property (63) satisfied by the exact PDE solution. Now split the first of the two remaining integrals over $x\in[\widetilde{\xi}_{i+1/2},\widetilde{x}_{i+1/2}]$ and $x\in[\widetilde{x}_{i+1/2},x_{i+1/2}]$; next, recall from discussion in Section SM3.1 that the integral over $[\widetilde{x}_{i+1/2},x_{i+1/2}]$ will be zero (it is an integral over an integer number of FV cells). Carrying out the analogous steps for the second integral and then invoking the definition of the error $\widetilde{E}_{i+1/2}^{n}$ from (SM14) gives the claim.
Lemma SM3.3 (Error estimate for semi-Lagrangian numerical flux).
Suppose the assumptions of Lemma SM3.1 hold, and that the numerical flux is evaluated as described in Section SM3.1. Further suppose that the solution $e(x,t)$ of (cons-lin) is sufficiently smooth, and that the wave-speed function in (cons-lin) varies only in time, $\alpha(x,t)\equiv\alpha(t)$. Then, the error (SM14) of the numerical flux satisfies
$\widetilde{E}_{i+1/2}^{n}=-h^{p+1}g_{p+1}(\varepsilon_{i+1/2})\frac{\partial^{p}e}{\partial x^{p}}\bigg{|}_{(x_{i+1/2},t_{n+1})}+{\cal O}(h^{p+2}),$ (SM17)
@@
$g(z):=\frac{1}{(p+1)}\prod\limits_{j=-k}^{k-1}(z+j).$ (SM18)
Proof SM3.4.
For notational simplicity let us momentarily drop temporal dependence. To begin, recall from Section SM3.1 that $R_{i+1/2}[\bar{\bm{e}}](x)\equiv Q_{p}^{\prime}(x)$, with $Q_{p}(x)$ a polynomial interpolating $E(x)$, the primitive of $e(x)$, i.e., $e(x)=E^{\prime}(x)$, at the $p+1$ cell interfaces in an interpolation stencil centered about the departure point $\widehat{\xi}_{i+1/2}$. Thus, by the fundamental theorem of calculus, the flux error (SM14) can be expressed as
$\widetilde{E}_{i+1/2}$ $:=\int_{\widetilde{\xi}_{i+1/2}}^{\widetilde{x}_{i+1/2}}R_{i+1/2}[\bar{\bm{e}}](x)-e(x)\operatorname{d}\!x,$ (SM19)
$=\big{[}Q_{p}(\widetilde{\xi}_{i+1/2})-E(\widetilde{\xi}_{i+1/2})\big{]}-\big{[}Q_{p}(\widetilde{x}_{i+1/2})-E(\widetilde{x}_{i+1/2})\big{]},$ (SM20)
$=Q_{p}(\widetilde{\xi}_{i+1/2})-E(\widetilde{\xi}_{i+1/2}),$ (SM21)
@@
Now to estimate (SM21) we use standard polynomial interpolation theory in the sense of Lemma SM1.1. To this end, we first restate the result of Lemma SM1.1 as it applies to the current context, i.e., with $U\mapsto E$, $k\mapsto 2k-1=p$ and $\ell=k-1$. Specifically, for any $x\in[\widetilde{x}_{i+1/2}-kh,\widetilde{x}_{i+1/2}+(k-1)h]$, we have
$E(x)-Q_{p}(x)=\frac{1}{(p+1)!}\prod\limits_{j=-k}^{k-1}\big{[}x-(\widetilde{x}_{i+1/2}+jh)\big{]}E^{(p+1)}(\varphi(x)),$ (SM22)
for $\varphi(x)\in[\widetilde{x}_{i+1/2}-kh,\widetilde{x}_{i+1/2}+(k-1)h]$ an unknown point. Now evaluate this error estimate at $x=\widetilde{x}_{i+1/2}-h\varepsilon_{i+1/2}$. First, let us reintroduce temporal dependence since it is now important. Noting that $\widetilde{x}_{i+1/2}-h\varepsilon_{i+1/2}-(\widetilde{x}_{i+1/2}+jh)=-h(\varepsilon_{i+1/2}+j)$ and that $p$ is odd, we find from (SM21)
$\widetilde{E}_{i+1/2}^{n}$ $=-\frac{h^{p+1}}{(p+1)!}\prod\limits_{j=-k}^{k-1}(\varepsilon_{i+1/2}+j)\frac{\partial^{p+1}E}{\partial x^{p+1}}\bigg{|}_{(\varphi(\widetilde{\xi}_{i+1/2}),t_{n})},$ (SM23)
$=-h^{p+1}g_{p+1}(\varepsilon_{i+1/2})\frac{\partial^{p}e}{\partial x^{p}}\bigg{|}_{(\varphi(\widetilde{\xi}_{i+1/2}),t_{n})},$ (SM24)
with $g_{p+1}$ defined in (SM18).
Next, note that since $\varphi(\widetilde{\xi}_{i+1/2})\in[\widetilde{x}_{i+1/2}-kh,\widetilde{x}_{i+1/2}+(k-1)h]$, we have $\varphi(\widetilde{\xi}_{i+1/2})=\widetilde{\xi}_{i+1/2}+{\cal O}(h)$ because $\widetilde{\xi}_{i+1/2}\in[\widetilde{x}_{i+1/2}-h,\widetilde{x}_{i+1/2}]$. Therefore, by Taylor expansion, $\frac{\partial^{p}e}{\partial x^{p}}\big{|}_{(\varphi(\widetilde{\xi}_{i+1/2}),t_{n})}=\frac{\partial^{p}e}{\partial x^{p}}\big{|}_{(\widetilde{\xi}_{i+1/2},t_{n})}+{\cal O}(h)$. Finally, we note that differentiating the PDE in question, $\frac{\partial}{\partial t}(e)+\frac{\partial}{\partial x}(\alpha(t)e)=0$, $p$ times with respect to $x$ gives $\frac{\partial}{\partial t}\big{(}\frac{\partial^{p}e}{\partial x^{p}}\big{)}+\frac{\partial}{\partial x}\big{(}\alpha(t)\frac{\partial^{p}e}{\partial x^{p}}\big{)}=0$. Thus, the derivative $\frac{\partial^{p}e}{\partial x^{p}}$ is constant along the same characteristics as the PDE in question, and, so, $\frac{\partial^{p}e}{\partial x^{p}}\big{|}_{(\widetilde{\xi}_{i+1/2},t_{n})}=\frac{\partial^{p}e}{\partial x^{p}}\big{|}_{(x_{i+1/2},t_{n+1})}$ because $(x_{i+1/2},t_{n+1})$ is the arrival point associated with the departure point $(\widetilde{\xi}_{i+1/2},t_{n})$; see Fig. 1. Inserting this into (SM24) concludes the proof.
SM3.3 Approximating truncation error
@@
Combining the results of Lemmas SM3.1 and SM3.3, we suppose that the truncation error for a single time-step of the coarse-grid semi-Lagrangian discretization method used in (58) satisfies
$\tau^{n}_{i}$ $:=\Big{(}\bar{\bm{e}}(t_{n+1})-{\cal S}^{t_{n},m\delta t}_{p,1_{*}}\bar{\bm{e}}(t_{n})\Big{)}_{i}$ (SM25)
$\approx-h^{p+1}\frac{\displaystyle{g_{p+1}(\varepsilon_{i+1/2})\frac{\partial^{p}e}{\partial x^{p}}\bigg{|}_{(x_{i+1/2},t_{n+1})}-g_{p+1}(\varepsilon_{i-1/2})\frac{\partial^{p}e}{\partial x^{p}}\bigg{|}_{(x_{i-1/2},t_{n+1})}}}{h},$ (SM26)
$\approx-h^{p+1}\Big{(}{\cal D}_{1}\operatorname{diag}\big{(}g_{p+1}(\bm{\varepsilon})\big{)}{\cal D}_{p}^{\top}\bar{\bm{e}}(t_{n+1})\Big{)}_{i}.$ (SM27)
@@
Here, $\bm{\varepsilon}=\big{(}\varepsilon_{1+1/2},\ldots,\varepsilon_{n_{x}+1/2}\big{)}\in\mathbb{R}^{n_{x}}$, and the polynomial $g_{p+1}$ from (SM18) is applied elementwise. As previously, ${\cal D}_{j}\in\mathbb{R}^{n_{x}\times n_{x}}$ for $j$ odd is a 1st-order accurate FD approximation to the $j$th derivative of a periodic grid function, on a stencil that has a 1-point bias to the left.
@@
The second approximation (SM27) arises simply from discretizing the $p$th-order derivatives in (SM26). On the other hand, (SM26) arises from a series of approximations that are more difficult to justify. That being said, they are ultimately justified by the fact that they lead to an expression for ${\cal T}_{\rm direct}^{n}$ that results in fast MGRIT convergence. First, the error estimate Lemma SM3.1 assumed departure points are located exactly; however, in practice they are located approximately. Second, the error estimate Lemma SM3.3 applies only to spatially independent wave-speed functions $\alpha$; however, in practice the wave-speed function varies in space. Third, the higher-order terms from Lemma SM3.3 are assumed not to be important; however, without further analysis it is not clear that do not result in terms of the same order as those that have been retained.
Based on (SM27) and the above discussion, we implement the truncation error operator as
${\cal T}_{\rm direct}^{n}=-h^{p+1}{\cal D}_{1}\operatorname{diag}\big{(}g_{p+1}(\bm{\varepsilon})\big{)}{\cal D}_{p}^{\top}.$ (SM28)
SM4 Error estimates for linear(ized) method-of-lines discretizations
@@
In this section, we present an error analysis for a linear method-of-lines discretization for a PDE that is an approximation to the linear conservation law (cons-lin). The outcome of this analysis is that it is used to develop an expression for the approximate truncation error operator ${\cal T}_{\rm ideal}^{n}$ in the MGRIT coarse-grid operator (58) that is used to solve certain linearized discretizations of (cons-lin) (the development of ${\cal T}_{\rm ideal}^{n}$ is given later in Section SM6). In addition, the analysis in this section is used to develop an MGRIT coarse-grid operator for solving standard, linear method-of-lines discretizations of (cons-lin) (these developments come next in Section SM5).
@@
The remainder of this section is structured as follows. Section SM4.1 presents a simplified model PDE and describes its discretization. Section SM4.2 presents a spatial-discretization-agnostic truncation error estimate for the model problem and discretization. Section SM4.3 presents error estimates for some standard linear reconstructions.
SM4.1 A simplified linear PDE and its discretization
@@
While it is ultimately discretizations of the PDE (cons-lin) that we solve with MGRIT, we do not develop truncation error estimates for this equation directly, but instead for a semi-autonomous approximation to it. [^8] In essence, two key components in our analysis are computing the exact solution of a linear ODE system of the form $\operatorname{d}\!\bar{\bm{e}}(t)/\operatorname{d}\!t=L\bar{\bm{e}}(t)$, and expressing in closed-form the local truncation error of its Runge-Kutta approximation. Both of these tasks are straightforward only when $L$ is time independent. Aspects of the analysis also apply to standard, linear method-of-lines discretizations of the PDE (cons-lin), enabling the development of an MGRIT solver for such problems; this scenario is considered in Section SM5.

### v2

@@ Acknowledgments
In this section, we consider further details of the conservative, FV semi-Lagrangian method outlined in Section 5.3. Section SM5.1 presents relevant details for how the numerical flux is implemented, Section SM5.2 develops a truncation error estimate, and then Section SM5.3 describes how this estimate is used to develop the expression for the associated truncation error operator ${\cal T}_{\rm direct}^{n}$ used in the coarse-grid operator (40).
SM5.1 Further details on numerical flux
@@
(SM22) $\widetilde{f}_{i+1/2}=\int_{\widetilde{\xi}_{i+1/2}}^{\widetilde{x}_{i+1/2}}R_{i+1/2}[\bar{\bm{e}}^{n}](x)\operatorname{d}\!x+\int_{\widetilde{x}_{i+1/2}}^{x_{i+1/2}}R_{i+1/2}[\bar{\bm{e}}^{n}](x)\operatorname{d}\!x,$
where, recall, $\widetilde{x}_{i+1/2}$ is the FV cell interface immediately to the east of the departure point $\widetilde{\xi}_{i+1/2}$; see Figure 1. The second integral in (SM22) is evaluated exactly, noting that it is simply a sum of cell averages across the cells in the interval $x\in[\widetilde{x}_{i+1/2},x_{i+1/2}]$. The first integral in (SM22) is more complicated. Recall that $R_{i+1/2}[\bar{\bm{e}}^{n}](x)$ is a polynomial reconstruction of $e^{n}(x)$ and it is constructed by enforcing that it preserves cell averages of $e^{n}(x)$ across the FV cells used in its stencil. For the remainder of this discussion, let us drop temporal subscripts since they are not important.
@@
Recall from Section 5.3 that we consider semi-Lagrangian methods of order p=2k1p=2k-1 in space, with $k=1,2,\ldots$. To this end, let $R_{i+1/2}[\bar{\bm{e}}^{n}](x)$ be a reconstruction polynomial using a stencil consisting of p=2k1p=2k-1 cells. Further, suppose the stencil is centered about the cell containing the departure point. For simplicity of notation, suppose that the cell containing the departure point, i.e., $x\in[\widetilde{x}_{i+1/2}-h,\widetilde{x}_{i+1/2}]$, is cell ${\cal I}_{0}$. Then, the cells in the stencil are $\{{\cal I}_{-(k-1)},\ldots,{\cal I}_{k-1}\}$. [^7] In the notation of Section SM1, the stencil has a total number of cells $k\mapsto 2k-1=p$, and a left shift of $\ell=k-1$. Define $E(x)$ as the primitive of $e(x)$, i.e., $E(x)=\int_{-\infty}^{x}e(\zeta)\operatorname{d}\!\zeta$ and $E^{\prime}(x)=e(x)$. Then, based on Section SM1, we know that $R_{i+1/2}[\bar{\bm{e}}](x)$ is a polynomial of degree at most p1p-1, and that it is the derivative of an interpolating polynomial $Q_{p}(x)$ of degree at most $p$,
(SM23) $R_{i+1/2}[\bar{\bm{e}}](x)\equiv Q_{p}^{\prime}(x),$
@@
(SM24) $Q_{p}(x_{j\pm 1/2})=E(x_{j\pm 1/2}),\quad j=-(k-1),\ldots,k-1.$
SM5.2 Error estimate
The estimate we develop herein and its proof share close parallels with the error estimate we developed previously in [14, Lemma 3.1] for a non-conservative, FD semi-Lagrangian method; however, the estimate itself and its proof are fundamentally different due to the fact that we are now considering a semi-Lagrangian method that is both conservative and based on finite volumes. The full estimate is split across Lemmas SM5.1 and SM5.3. Note that the proof of Lemma SM5.3 in particular requires to have read Section SM5.1.
Lemma SM5.1 (Error estimate for semi-Lagrangian method).
@@
(SM25) $\tau^{n}_{i}$ $:=\Big(\bar{\bm{e}}(t_{n+1})-{\cal S}^{t_{n},\delta t}_{p,\infty}\bar{\bm{e}}(t_{n})\Big)_{i}=\frac{\widetilde{E}_{i+1/2}^{n}-\widetilde{E}_{i-1/2}^{n}}{h},$
@@
(SM26) $\widetilde{E}_{i+1/2}^{n}$ $:=\int_{\widetilde{\xi}_{i+1/2}}^{\widetilde{x}_{i+1/2}}R_{i+1/2}[\bar{\bm{e}}(t_{n})](x)-e(x,t_{n})\operatorname{d}\!x.$
Proof SM5.2.
@@
(SM27) $\tau^{n}_{i}$ $=\bar{e}_{i}(t_{n+1})-\bar{e}_{i}(t_{n})+\frac{\int_{\widetilde{\xi}_{i+1/2}}^{x_{i+1/2}}R_{i+1/2}[\bar{\bm{e}}(t_{n})](x)\operatorname{d}\!x-\int_{\widetilde{\xi}_{i-1/2}}^{x_{i-1/2}}R_{i-1/2}[\bar{\bm{e}}(t_{n})](x)\operatorname{d}\!x}{h}.$
@@
(SM28) $\begin{split}&\tau^{n}_{i}=\left[\bar{e}_{i}(t_{n+1})-\frac{\int_{\widetilde{\xi}_{i-1/2}}^{x_{i-1/2}}e(x,t_{n})\operatorname{d}\!x+\int_{x_{i-1/2}}^{x_{i+1/2}}e(x,t_{n})\operatorname{d}\!x+\int_{x_{i+1/2}}^{\widetilde{\xi}_{i+1/2}}e(x,t_{n})\operatorname{d}\!x}{h}\right]+\\ &\frac{\Big(\int_{\widetilde{\xi}_{i+1/2}}^{x_{i+1/2}}R_{i+1/2}[\bar{\bm{e}}(t_{n})](x)-e(x,t_{n})\operatorname{d}\!x\Big)-\Big(\int_{\widetilde{\xi}_{i-1/2}}^{x_{i-1/2}}R_{i-1/2}[\bar{\bm{e}}(t_{n})](x)-e(x,t_{n})\operatorname{d}\!x\Big)}{h}.\end{split}$
The term inside the closed parentheses vanishes due to the local conservation property (45) satisfied by the exact PDE solution. Now split the first of the two remaining integrals over $x\in[\widetilde{\xi}_{i+1/2},\widetilde{x}_{i+1/2}]$ and $x\in[\widetilde{x}_{i+1/2},x_{i+1/2}]$; next, recall from discussion in Section SM5.1 that the integral over $\widetilde{x}_{i+1/2},x_{i+1/2}]$ will be zero (it is an integral over an integer number of FV cells). Carrying out the analogous steps for the second integral and then invoking the definition of the error $\widetilde{E}_{i+1/2}^{n}$ from (SM26) gives the claim.
Lemma SM5.3 (Error estimate for semi-Lagrangian numerical flux).
Suppose the assumptions of Lemma SM5.1 hold, and that the numerical flux is evaluated as described in Section SM5.1. Further suppose that the solution $e(x,t)$ of (cons-lin) is sufficiently smooth, and that the wave-speed function in (cons-lin) varies only in time, $\alpha(x,t)\equiv\alpha(t)$. Then, the error (SM26) of the numerical flux satisfies
(SM29) $\widetilde{E}_{i+1/2}^{n}=-h^{p+1}g_{p+1}(\varepsilon_{i+1/2})\frac{\partial^{p}e}{\partial x^{p}}\bigg|_{(x_{i+1/2},t_{n+1})}+{\cal O}(h^{p+2}),$
@@
(SM30) $g(z):=\frac{1}{(p+1)}\prod\limits_{j=-k}^{k-1}(z+j).$
Proof SM5.4.
For notational simplicity let us momentarily drop temporal dependence. To begin, recall from Section SM5.1 that $R_{i+1/2}[\bar{\bm{e}}](x)\equiv Q_{p}^{\prime}(x)$, with $Q_{p}(x)$ a polynomial interpolating $E(x)$, the primitive of $e(x)$, i.e., $e(x)=E^{\prime}(x)$, at the $p+1$ cell interfaces in an interpolation stencil centered about the departure point $\widehat{\xi}_{i+1/2}$. Thus, by the fundamental theorem of calculus, the flux error (SM26) can be expressed as
(SM31) $\widetilde{E}_{i+1/2}$ $:=\int_{\widetilde{\xi}_{i+1/2}}^{\widetilde{x}_{i+1/2}}R_{i+1/2}[\bar{\bm{e}}](x)-e(x)\operatorname{d}\!x,$
(SM32) $=\big[Q_{p}(\widetilde{\xi}_{i+1/2})-E(\widetilde{\xi}_{i+1/2})\big]-\big[Q_{p}(\widetilde{x}_{i+1/2})-E(\widetilde{x}_{i+1/2})\big],$
(SM33) $=Q_{p}(\widetilde{\xi}_{i+1/2})-E(\widetilde{\xi}_{i+1/2}),$
@@
Now to estimate (SM33) we use standard polynomial interpolation theory in the sense of Lemma SM1.1. To this end, we first restate the result of Lemma SM1.1 as it applies to the current context, i.e., with $U\mapsto E$, $k\mapsto 2k-1=p$ and $\ell=k-1$. Specifically, for any $x\in[\widetilde{x}_{i+1/2}-kh,\widetilde{x}_{i+1/2}+(k-1)h]$, we have
(SM34) $E(x)-Q_{p}(x)=\frac{1}{(p+1)!}\prod\limits_{j=-k}^{k-1}\big[x-(\widetilde{x}_{i+1/2}+jh)\big]E^{(p+1)}(\varphi(x)),$
for $\varphi(x)\in[\widetilde{x}_{i+1/2}-kh,\widetilde{x}_{i+1/2}+(k-1)h]$ an unknown point. Now evaluate this error estimate at $x=\widetilde{x}_{i+1/2}-h\varepsilon_{i+1/2}$. First, let us reintroduce temporal dependence since it is now important. Noting that $\widetilde{x}_{i+1/2}-h\varepsilon_{i+1/2}-(\widetilde{x}_{i+1/2}+jh)=-h(\varepsilon_{i+1/2}+j)$ and that $p$ is odd, we find from (SM33)
(SM35) $\widetilde{E}_{i+1/2}^{n}$ $=-\frac{h^{p+1}}{(p+1)!}\prod\limits_{j=-k}^{k-1}(\varepsilon_{i+1/2}+j)\frac{\partial^{p+1}E}{\partial x^{p+1}}\bigg|_{(\varphi(\widetilde{\xi}_{i+1/2}),t_{n})},$
(SM36) $=-h^{p+1}g_{p+1}(\varepsilon_{i+1/2})\frac{\partial^{p}e}{\partial x^{p}}\bigg|_{(\varphi(\widetilde{\xi}_{i+1/2}),t_{n})},$
with $g_{p+1}$ defined in (SM30).
Next, note that since $\varphi(\widetilde{\xi}_{i+1/2})\in[\widetilde{x}_{i+1/2}-kh,\widetilde{x}_{i+1/2}+(k-1)h]$, we have $\varphi(\widetilde{\xi}_{i+1/2})=\widetilde{\xi}_{i+1/2}+{\cal O}(h)$ because $\widetilde{\xi}_{i+1/2}\in[\widetilde{x}_{i+1/2}-h,\widetilde{x}_{i+1/2}]$. Therefore, by Taylor expansion, $\frac{\partial^{p}e}{\partial x^{p}}\big|_{(\varphi(\widetilde{\xi}_{i+1/2}),t_{n})}=\frac{\partial^{p}e}{\partial x^{p}}\big|_{(\widetilde{\xi}_{i+1/2},t_{n})}+{\cal O}(h)$. Finally, we note that differentiating the PDE in question, $\frac{\partial}{\partial t}(e)+\frac{\partial}{\partial x}(\alpha(t)e)=0$, $p$ times with respect to $x$ gives $\frac{\partial}{\partial t}\big(\frac{\partial^{p}e}{\partial x^{p}}\big)+\frac{\partial}{\partial x}\big(\alpha(t)\frac{\partial^{p}e}{\partial x^{p}}\big)=0$. Thus, the derivative $\frac{\partial^{p}e}{\partial x^{p}}$ is constant along the same characteristics as the PDE in question, and, so, $\frac{\partial^{p}e}{\partial x^{p}}\big|_{(\widetilde{\xi}_{i+1/2},t_{n})}=\frac{\partial^{p}e}{\partial x^{p}}\big|_{(x_{i+1/2},t_{n+1})}$ because $(x_{i+1/2},t_{n+1})$ is the arrival point associated with the departure point $(\widetilde{\xi}_{i+1/2},t_{n})$; see Figure 1. Inserting this into (SM36) concludes the proof.
SM5.3 Approximating truncation error
@@
Combining the results of Lemmas SM5.1 and SM5.3, we suppose that the truncation error for a single time-step of the coarse-grid semi-Lagrangian discretization method used in (40) satisfies
(SM37) $\tau^{n}_{i}$ $:=\Big(\bar{\bm{e}}(t_{n+1})-{\cal S}^{t_{n},m\delta t}_{p,1_{*}}\bar{\bm{e}}(t_{n})\Big)_{i}$
(SM38) $\approx-h^{p+1}\frac{\displaystyle{g_{p+1}(\varepsilon_{i+1/2})\frac{\partial^{p}e}{\partial x^{p}}\bigg|_{(x_{i+1/2},t_{n+1})}-g_{p+1}(\varepsilon_{i-1/2})\frac{\partial^{p}e}{\partial x^{p}}\bigg|_{(x_{i-1/2},t_{n+1})}}}{h},$
(SM39) $\approx-h^{p+1}\Big({\cal D}_{1}\operatorname{diag}\big(g_{p+1}(\bm{\varepsilon})\big){\cal D}_{p}^{\top}\bar{\bm{e}}(t_{n+1})\Big)_{i}.$
@@
Here, $\bm{\varepsilon}=\big(\varepsilon_{1+1/2},\ldots,\varepsilon_{n_{x}+1/2}\big)\in\mathbb{R}^{n_{x}}$, and the polynomial $g_{p+1}$ from (SM30) is applied elementwise. As previously, ${\cal D}_{j}\in\mathbb{R}^{n_{x}\times n_{x}}$ for $j$ odd is a 1st-order accurate FD approximation to the $j$th derivative of a periodic grid function, on a stencil that has a 1-point bias to the left.
@@
The second approximation (SM39) arises simply from discretizing the $p$th-order derivatives in (SM38). On the other hand, (SM38) arises from a series of approximations that are more difficult to justify. That being said, they are ultimately justified by the fact that they lead to an expression for ${\cal T}_{\rm direct}^{n}$ that results in fast MGRIT convergence. First, the error estimate Lemma SM5.1 assumed departure points are located exactly; however, in practice they are located approximately. Second, the error estimate Lemma SM5.3 applies only to spatially independent wave-speed functions $\alpha$; however, in practice the wave-speed function varies in space. Third, the higher-order terms from Lemma SM5.3 are assumed not to be important; however, without further analysis it is not clear that do not result in terms of the same order as those that have been retained.
Based on (SM39) and the above discussion, we implement the truncation error operator as
(SM40) ${\cal T}_{\rm direct}^{n}=-h^{p+1}{\cal D}_{1}\operatorname{diag}\big(g_{p+1}(\bm{\varepsilon})\big){\cal D}_{p}^{\top}.$
SM6 Error estimates for linear(ized) method-of-lines discretizations
@@
In this section, we present an error analysis for a linear method-of-lines discretization for a PDE that is an approximation to the linear conservation law (cons-lin). The outcome of this analysis is that it is used to develop an expression for the approximate truncation error operator ${\cal T}_{\rm ideal}^{n}$ in the MGRIT coarse-grid operator (40) that is used to solve certain linearized discretizations of (cons-lin) (the development of ${\cal T}_{\rm ideal}^{n}$ is given later in Section SM8). In addition, the analysis in this section is used to develop an MGRIT coarse-grid operator for solving standard, linear method-of-lines discretizations of (cons-lin) (these developments come next in Section SM7).
@@
The remainder of this section is structured as follows. Section SM6.1 presents a simplified model PDE and describes its discretization. Section SM6.2 presents a spatial-discretization-agnostic truncation error estimate for the model problem and discretization. Section SM6.3 presents error estimates for some standard linear reconstructions.
SM6.1 A simplified linear PDE and its discretization
@@
While it is ultimately discretizations of the PDE (cons-lin) that we solve with MGRIT, we do not develop truncation error estimates for this equation directly, but instead for a semi-autonomous approximation to it. [^8] In essence, two key components in our analysis are computing the exact solution of a linear ODE system of the form $\operatorname{d}\!\bar{\bm{e}}(t)/\operatorname{d}\!t=L\bar{\bm{e}}(t)$, and expressing in closed-form the local truncation error of its Runge-Kutta approximation. Both of these tasks are straightforward only when $L$ is time independent. Aspects of the analysis also apply to standard, linear method-of-lines discretizations of the PDE (cons-lin), enabling the development of an MGRIT solver for such problems; this scenario is considered in Section SM7.